Farmer Jim is building a new chicken coop in the shape of a rectangular prism. He needs the height of the coop to be 2 feet shorter than the width, and needs the length of the coop to be 10 feet longer than the width.
Create an equation to represent the volume of the chicken coop, y, in terms of the width, x, in feet. Enter the equation in the field by replacing the values of a, b, c, and d.

Answer by Boreal(15235)

width is x
height is x-2
length is x+10
volume is the product of all of these
(x^2-2x)(x+10)=x^3+10x^2-2x^2-20x
y=x^3+8x^2-20x